# MobiFuse: Learning Universal Human Mobility Patterns through Cross-domain Data Fusion

### v1

@@ C.1.2 Base Model and Its Transferability
Essential activities such as home, work, and school show closely aligned patterns between the two datasets, evidenced by low JSD values (0.023, 0.021, and 0.024 respectively), highlighting the universality of these routines. Conversely, regional variations are apparent in other activities. Shopping patterns, for example, differ significantly, with SCAG data showing a midday peak (JSD 0.054) versus NHTS’s more even distribution, suggesting distinct regional shopping behaviors. Exercise activities also vary, with NHTS indicating a preference for early morning sessions (JSD 0.036). These differences likely mirror varying lifestyle choices between urban Los Angeles and the broader national context. The model’s consistent low $M^{S}$ and $M^{N}$ values, typically below 0.01, confirm its efficacy in capturing these nuances, underscoring its utility for urban planning and policy-making by identifying both universal and regional behavioral patterns.
The JSD values for the evaluation metrics across iterations are presented in Table 3. Only three metrics are evaluated, because activity duration and activity chain length lack ground truth in incomplete data, making them unreliable for comparison. In contrast, the distributions of activity type, start time, and end time are more accurate at the system level and effectively reflect regional mobility patterns. The Table 3 shows the iterative transfer learning process improving alignment with the target region. Activity type JSD decreases significantly from 0.0508 (baseline) to a minimum of 0.0022 at Iteration 3, demonstrating effective adaptation. Start time and end time JSDs reach their lowest values (0.0007 and 0.0019, respectively) at Iteration 4. However, slight increases in later iterations suggest diminishing returns or potential overfitting. Overall, the results highlight that Iterations 3 and 4 are optimal for reducing dissimilarities across these metrics, reflecting successful model adaptation. Figure 9 demonstrates how the progressive adaptation of the NHTS-based model to the LA context over multiple iterations via the proposed Semi-Supervised Transfer Learning using LA GPS data.
@@
The second step of transfer learning is performing the demographic adaptation by fusing TUS data. We trained 12 demographic adapters in accordance with sex (male/female), age (young, mid-aged, elderly), and employment status (worker, non-worker). To establish target distributions for each demographic group, we aggregate TUS data to create activity pattern benchmarks. Analysis of the aggregated Time Use Survey distributions (Table 5) reveals distinct activity patterns across demographic groups with aspects of sex, age, and employment status. We proceed with adapter training for each demographic group. The training process demonstrates varying levels of convergence and efficiency across different demographic segments. Overall, the adaptation process shows strong performance, with reward values (measuring similarity to target distributions) consistently above 0.85 across all demographic groups, and many groups achieving rewards above 0.95.
@@
Young Male W. (0.0085) 843.71/837.50 209.58/224.50 117.76/134.74 19.61/15.01 52.49/38.51 100.71/53.32 23.82/18.55 61.53/53.38
Young Female W. (0.0159) 820.18/785.27 176.23/204.68 131.50/134.76 39.82/20.84 57.31/34.26 70.19/34.58 23.50/18.79 75.84/88.82
Mid-Aged Male W. (0.0189) 798.36/800.37 380.79/382.87 0.00/0.00 37.89/19.10 61.70/76.96 76.16/62.16 23.92/4.67 42.35/0.98
Mid-Aged Female W. (0.0185) 839.04/825.56 332.97/330.77 0.00/0.00 42.36/22.70 57.00/74.31 63.22/58.69 16.75/4.43 46.14/2.07
Elderly Male W. (0.0157) 875.00/886.00 295.59/284.50 0.00/0.00 34.75/29.95 69.24/77.71 83.23/61.21 20.51/4.41 35.58/0.80
Elderly Female W. (0.0136) 913.23/902.44 245.65/260.23 0.00/0.00 41.41/35.62 62.34/67.56 84.74/60.80 15.03/3.99 36.99/0.83
Young Male N. (0.0090) 903.82/960.43 0.00/0.00 170.74/203.06 21.34/14.47 54.04/68.70 97.69/84.56 57.86/41.07 52.06/44.74
Young Female N. (0.0103) 899.88/987.70 0.00/0.00 170.54/149.18 36.01/21.77 59.85/75.15 79.72/76.87 31.51/39.91 49.20/40.00
Mid-Aged Male N. (0.0064) 1013.83/1054.20 0.00/0.00 18.49/3.80 40.56/42.97 57.17/69.00 127.28/118.24 26.97/37.15 45.55/32.73
Mid-Aged Female N. (0.0078) 993.96/1083.50 0.00/0.00 11.88/3.86 44.16/22.48 58.29/64.68 90.56/67.73 16.93/11.26 54.37/57.17
Elderly Male N. (0.0092) 1022.03/1101.41 0.00/0.00 1.56/3.34 42.11/43.75 70.30/80.47 144.36/150.71 25.04/5.76 41.62/22.40
Elderly Female N. (0.0099) 1037.10/1115.56 0.00/0.00 1.63/3.32 37.40/41.37 63.45/88.91 120.76/128.88 15.08/4.74 46.89/31.46
@@
Note: Values shown as TUS/Output. W: Worker, N: Non-worker. JSD values indicate distribution similarity (lower is better).
The entire adapter training process, implemented on an NVIDIA RTX 4090 GPU, requires approximately 10 hours of computation time, averaging 40 minutes per demographic group. This relatively efficient training time, despite the complexity of demographic-specific pattern learning, demonstrates the practical feasibility of our approach for real-world applications.
After training the demographic-aware adapters, we integrate them with the transferred reconstruction model to create a comprehensive system capable of generating both demographically appropriate and temporally coherent activity patterns. This integrated model preserves the base model’s capabilities in temporal pattern recognition and activity dependencies while incorporating demographic sensitivity through the trained adapters.
Using the trained demographic-aware model, we conduct large-scale inference to generate synthetic activity patterns. We handle different trajectory data using different masking methods. For trajectories with relatively low loss or high completeness, we can often infer their socio-demographic information. In such cases, the problem becomes one of completion: we simply match the corresponding socio-demographic info to the trajectory and use the model to fill in the missing data. However, for data with low completeness, it is not possible to infer any socio-demographic information. We treat this as a generation task, masking as much data as possible—up to 80% of the time in a day—and assigning it a socio-demographic profile. The model then generates a trajectory consistent with this socio-demographic information.
Figure 10: Comparison of aggregated activity start time distributions across data sources (NHTS Survey, Veraset GPS Data, and Model-Generated Data)
The model generates activity chains through multiple passes - 5 passes for weekdays and 2 passes for weekends - with each pass using different random masks to ensure diversity in the generated patterns. This process allows the model to explore various possible activity combinations while maintaining demographic consistency. The generation process incorporates both demographic factors and variations in the day of the week to ensure realistic activity patterns between different segments of the population and throughout the week.
The generation demonstrates strong alignment with Time Use Survey across all demographic groups. Quantitative evaluation using JSD shows excellent performance (detailed results in Table 5), with non-worker groups achieving particularly strong alignment (JSD 0.0064-0.0103) and worker groups showing comparable performance (JSD 0.0085-0.0189). Mid-aged male non-workers achieve the best performance (JSD = 0.0064), while maintaining realistic activity patterns such as work duration differences across age groups and gender-specific activity preferences.
@@
To validate our transfer learning fusing module successfully learns the activity and demographic nuances from multi-source data, we analyze start time distributions across major activity types, comparing our generated patterns with both NHTS and raw GPS input data. The analysis is conducted at an aggregate level rather than per demographic group to evaluate the model’s overall capability in preserving temporal dynamics and to provide sufficient sample sizes for meaningful comparison, particularly for less frequent activities. We visualize and quantify these comparisons through probability distributions and JSD metrics to assess both pattern similarity and statistical significance.
As shown in Figure 10, the aggregated activity start time distribution demonstrates strong alignment across all other data sources. The generated patterns show particularly close agreement with the raw GPS data (JSD: 0.004), confirming the model’s ability to preserve real-world temporal patterns while incorporating demographic variations. This corresponds to findings from (Stopher & Greaves, 2007), which highlight the temporal precision of GPS data compared to traditional survey methods. The comparison with NHTS data also shows strong correspondence (JSD: 0.014), with matching morning (8:00–10:00) and afternoon (16:00–18:00) peak periods, though with slight probability differences of approximately 0.005. These discrepancies reflect the systematic differences between survey and GPS data collection methods, such as recall bias and rounding effects, where respondents often report activities at common clock hours (Bricka et al., 2012), while GPS data captures actual movement patterns with higher temporal precision. At the individual activity level, as shown in Figure 11, the model maintains high fidelity across both major mandatory and non-mandatory activities, with average JSD values consistently below 0.06. We noticed that the generation distribution shapes strike a balance between the disagreement of GPS and Survey data and create a reasonable temporal activity patterns credit to the robust learning capabilities of our fusion and transfer learning pipeline.
C.1.3 Spatial Level Validation
We utilize the generated output from prior pipeline to assign locations in the context of LA County using Algorithm 1, with a spatial division of eight sub-regions (Figure 12(a)). The model is first calibrated using 100,000 agents from the SCAG Activity-Based Model (ABM) dataset, then scaled to generate mobility patterns for 1 million agents.
@@ C.1.3 Spatial Level Validation
Regional-level Analysis: The spatial distribution of activities across sub-regions demonstrates the effectiveness of our location assignment method (Figure 12(a)). The high cosine similarity (0.997) between the origin-destination matrices of our model and SCAG ABM validates the preservation of regional flow patterns.
@@
System-wide Traffic Performance: Traffic simulation on the LA transportation network (Figure 12(b)) reveals strong system-level performance. The hourly VMT and traffic speed patterns closely match the benchmark SCAG ABM results, with MAPEs of 4.97 and 1.16 respectively (Figure 13(b)). These metrics indicate accurate reproduction of aggregate travel patterns.
@@
Corridor-level Analysis: We conducted detailed analysis on a major segment of Interstate 405, comparing our results with both SCAG ABM predictions and PeMS real-world observations. The analysis reveals:
  • Directional variations in congestion patterns are well-captured
  • Northbound traffic: MAPE of 5.85 for volume, 4.45 for speed
  • Southbound traffic: MAPE of 9.32 for volume, 4.36 for speed
The corridor-level results (Figure 14) demonstrate the model’s capability to reproduce fine-grained traffic dynamics, including peak-hour congestion and directional flow patterns.
C.2 Egypt Case Study
C.2.1 Regional Adaptation in Data-Scarce Contexts
@@
C.2.2 Transfer Learning and Model Training
@@
Iteration_1 0.0055 0.0052 0.0027
Iteration_2 0.0038 0.0035 0.0021
Iteration_3 0.0039 0.0025 0.0021
Iteration_4 0.0027 0.0034 0.0029
Iteration_5 0.0043 0.0042 0.0045
Iteration_6 0.0029 0.0049 0.0046
To adapt the base model trained on the NHTS dataset, we employed transfer learning, leveraging the semi-supervised iterative refinement process detailed in our previous study. The adaptation process retained universal temporal patterns while progressively aligning with the distinct cultural and regional mobility characteristics of Egypt.
The progressive unfreezing approach was instrumental in balancing the preservation of global patterns with the learning of localized nuances. Initially, only the MLP and embedding layers were fine-tuned, allowing for rapid adaptation of activity classifications and temporal representations. In subsequent iterations, the transformer layers were gradually unfrozen to refine the model’s understanding of intermediate and complex activity dependencies. Its evolution through the semi-supervised transfer learning using Egypt GPS data is visualized and quantified in Figure 15 and Table 6.
@@
The Egypt TUS data was aggregated to derive target activity duration distributions across the six demographic groups. Adapter training followed the same computational configuration as the Los Angeles case, utilizing a single NVIDIA RTX 4090 GPU. Training each demographic group required approximately 40 minutes, demonstrating the scalability of the approach.
The inference process was conducted on NVIDIA L40S GPUs, generating synthetic activity chains that closely aligned with the target distributions. As shown in Table 7, the JSD values for all demographic groups were low (ranging from 0.0054 for mid-aged males to 0.0302 for mid-aged females), indicating strong alignment between the generated and target distributions. This underscores the framework’s ability to replicate realistic activity patterns even in a data-scarce context.
@@
Young Male (0.0067) 842.24/790.56 94.98/90.72 347.63/308.16 4.22/15.84 15.04/1.44 33.52/38.88 42.96/31.68 32.41/56.16
Young Female (0.0172) 786.10/761.76 76.75/70.56 316.58/313.92 3.13/53.28 14.41/4.32 131.45/148.32 1.43/8.64 36.24/53.28
Mid-Aged Male (0.0054) 684.13/688.32 342.37/348.48 12.70/10.08 8.87/14.40 15.00/15.84 136.65/126.72 4.07/4.32 119.86/112.32
Mid-Aged Female (0.0302) 881.72/832.32 199.05/146.88 63.73/0.00 8.07/24.48 19.69/5.76 146.69/126.72 0.33/4.32 36.20/53.28
Elderly Male (0.0141) 752.65/764.64 80.22/56.16 2.16/2.88 4.57/18.72 17.38/47.52 186.85/197.28 1.26/5.76 135.41/108.00
Elderly Female (0.0198) 848.55/853.92 107.94/79.20 0.00/17.28 3.10/18.72 10.85/4.32 164.29/142.56 0.00/11.52 46.59/106.56
@@
C.2.3 Observations on Generated Patterns
The generated activity chains demonstrated high consistency with target distributions, particularly for mandatory activities. Mid-aged males exhibited significantly longer work durations compared to mid-aged females, reflecting the disparity in employment rates (65% for males vs. 18% for females, as reported by the UN). Work durations for females in Egypt were notably lower than those in Los Angeles, illustrating the framework’s sensitivity to cultural and socio-economic differences. Religious Activities: Prominent across all demographic groups, with peaks reflecting the significance of prayer times in Egyptian society. Recreational Activities: Observed to be less frequent, consistent with the conservative cultural context of Egypt compared to Western regions like Los Angeles.

### v2

@@ 6.2.2 Base Model Evaluation on Survey Data
Essential activities such as home, work, and school show closely aligned patterns between the two datasets, evidenced by low JSD values in Table 3 (0.023, 0.021, and 0.024, respectively), highlighting the universality of these routines. Conversely, regional variations are apparent in other activities. Shopping patterns, for example, differ significantly, with SCAG data showing a midday peak (JSD 0.054) versus NHTS’s more even distribution, suggesting distinct regional shopping behaviors. Exercise activities also vary, with NHTS indicating a preference for early morning sessions (JSD 0.036). These differences likely mirror varying lifestyle choices between urban Los Angeles and the broader national context. The model’s consistently low $M^{S}$ and $M^{N}$ values, typically below 0.01, confirm its efficacy in capturing these nuances, underscoring its utility for urban planning and policy-making by identifying both universal and regional behavioral patterns.
JSD Values Length Duration Type Start Time End Time
NHTS vs SCAG 0.017 0.009 0.059 0.018 0.013
$M^{N}$ on NHTS 0.017 0.002 0.001 0.001 0.005
$M^{N}$ on SCAG 0.021 0.003 0.017 0.004 0.007
$M^{S}$ on SCAG 0.011 0.004 0.004 0.001 0.001
6.2.3 Transfer Learning Evaluation on GPS and Survey Data
JSD values across iterations for the evaluation metrics are shown in Table 4, showing how the iterative transfer learning process improves alignment with the target region. In this table, only three metrics are included because the lack of ground truth in incomplete data renders activity duration and chain length unreliable for comparison, whereas the selected metrics are more accurate at the system level and can better reflect regional mobility patterns. All values reported for Iteration 1 through Iteration 6 are the mean ± standard deviation over three random seeds. Activity type JSD drops from 0.0508 at baseline to 0.0028 by Iteration 3. Similarly, start time and end time JSDs reach their lowest values, 0.0008 and 0.0022, respectively, at Iteration 4. Model performance deteriorates beyond Iterations 3 and 4, which appear to strike the best balance for reducing dissimilarities and adapting the model. This decline aligns with findings that AI models collapse when trained on recursively generated data (Shumailov et al., 2024), as over-reliance on synthetic data erodes nuanced or rare activities in the original data. Figure 10 demonstrates how the progressive adaptation of the NHTS-based model to the LA context evolves over multiple iterations via the proposed Semi-Supervised Transfer Learning using LA GPS data.
Table 4: Model evolution during iterative transfer learning for LA case
JSD Values Activity Type Start Time End Time
NHTS vs LA GPS 0.0508 0.0289 0.0367
Iteration_1 0.0131 ± 0.0033 0.0018 ± 0.0006 0.0044 ± 0.0007
Iteration_2 0.0041 ± 0.0011 0.0015 ± 0.0004 0.0042 ± 0.0006
Iteration_3 0.0028 ± 0.0004 0.0010 ± 0.0002 0.0030 ± 0.0003
Iteration_4 0.0030 ± 0.0003 0.0008 ± 0.0001 0.0022 ± 0.0003
Iteration_5 0.0052 ± 0.0008 0.0027 ± 0.0005 0.0054 ± 0.0007
Iteration_6 0.0081 ± 0.0012 0.0050 ± 0.0009 0.0085 ± 0.0011
@@ 6.2.3 Transfer Learning Evaluation on GPS and Survey Data
The second step of transfer learning is performing the demographic adaptation by fusing TUS data. We trained 12 demographic adapters in accordance with sex (male/female), age (young, mid-aged, elderly), and employment status (worker, non-worker). Analysis of the aggregated Time Use Survey distributions (Table 6) reveals distinct activity patterns across demographic groups with aspects of sex, age, and employment status.
@@
Group HTS vs Output TUS vs Output TUS vs HTS
Young M .083/.086 .008/.009 .014/.014
Young F .094/.085 .015/.013 .016/.011
Mid-aged M .048/.067 .018/.006 .016/.054
Mid-aged F .055/.066 .018/.007 .022/.059
Elder M .070/.066 .015/.009 .020/.055
Elder F .077/.071 .013/.009 .029/.063
Note: Values shown as (Worker/Non-worker); M: Male F: Female
The generation demonstrates strong alignment with the Time Use Survey across all demographic groups. Quantitative evaluation using JSD shows excellent performance (detailed results in Table 6), with non-worker groups achieving particularly strong alignment (JSD 0.0064-0.0103) and worker groups showing comparable performance (JSD 0.0085-0.0189). Mid-aged male non-workers achieve the best performance (JSD = 0.0064) while maintaining realistic activity patterns, such as work duration differences across age groups and gender-specific activity preferences.
@@
Young Male W. (0.0085) 843.7/837.5 209.5/224.5 117.7/134.7 19.6/15.0 52.5/38.5 100.7/53.3 23.8/18.6 61.5/53.4
Young Female W. (0.0159) 820.2/785.3 176.2/204.7 131.5/134.8 39.8/20.8 57.3/34.3 70.2/34.6 23.5/18.8 75.8/88.8
Mid-Aged Male W. (0.0189) 798.4/800.4 380.8/382.9 0.0/0.0 37.9/19.1 61.7/77.0 76.2/62.2 23.9/4.7 42.4/1.0
Mid-Aged Female W. (0.0185) 839.0/825.6 333.0/330.8 0.0/0.0 42.4/22.7 57.0/74.3 63.2/58.7 16.8/4.4 46.1/2.1
Elderly Male W. (0.0157) 875.0/886.0 295.6/284.5 0.0/0.0 34.8/30.0 69.2/77.7 83.2/61.2 20.5/4.4 35.6/0.8
Elderly Female W. (0.0136) 913.2/902.4 245.7/260.2 0.0/0.0 41.4/35.6 62.3/67.6 84.7/60.8 15.0/4.0 37.0/0.8
Young Male N. (0.0090) 903.8/960.4 0.0/0.0 170.7/203.1 21.3/14.5 54.0/68.7 97.7/84.6 57.9/41.1 52.1/44.7
Young Female N. (0.0103) 899.9/987.7 0.0/0.0 170.5/149.2 36.0/21.8 59.9/75.2 79.7/76.9 31.5/39.9 49.2/40.0
Mid-Aged Male N. (0.0064) 1013.8/1054.2 0.0/0.0 18.5/3.8 40.6/43.0 57.2/69.0 127.3/118.2 27.0/37.2 45.6/32.7
Mid-Aged Female N. (0.0078) 994.0/1083.5 0.0/0.0 11.9/3.9 44.2/22.5 58.3/64.7 90.6/67.7 16.9/11.3 54.4/57.2
Elderly Male N. (0.0092) 1022.0/1101.4 0.0/0.0 1.6/3.3 42.1/43.8 70.3/80.5 144.4/150.7 25.0/5.8 41.6/22.4
Elderly Female N. (0.0099) 1037.1/1115.6 0.0/0.0 1.6/3.3 37.4/41.4 63.5/88.9 120.8/128.9 15.1/4.7 46.9/31.5
@@
Note: Values shown as TUS/Output. W: Worker, N: Non-worker. JSD values indicate distribution similarity (lower JSD means better result).
Beyond activity durations, we evaluate temporal patterns to ensure the demographic adapter preserves learned temporal dynamics while incorporating demographic variations. The temporal evaluation spans demographic groups (Figure 11) and activity-specific distributions (Figure 12). At the demographic level, mid-aged and elderly groups show consistent patterns across both worker and non-worker categories. The young group shows some discrepancies between TUS and HTS because the young worker group is insignificant in HTS (around 2% of the total population).
Figure 11: Temporal performance evaluation per demographic group based on. Note: HTS lacks a Young Worker group; comparison for young individuals ignores employment status.
At the aggregate level, the model maintains strong agreement with both raw GPS data (JSD: 0.004) and NHTS patterns (JSD: 0.014), capturing key features like morning (8:00-10:00) and afternoon (16:00-18:00) peaks. At the activity level, major activities like work and school are well captured and the patterns learned from the base model were maintained throughout the complex transfer learning process. Notably, our generated patterns and GPS data capture early morning activities (before 6 AM) that are absent in HTS data, addressing known survey limitations such as underreporting and rounding effects, for example, people typically fail to report some emergency travel activities and tend to make the pattern more routine (Bricka et al., 2012). This demonstrates our framework’s ability to balance GPS and survey data characteristics while maintaining demographic representativeness.
@@
At the activity level, as shown in Figure 12, major activities like work and school are well captured, and the patterns learned from the base model were maintained throughout the complex transfer learning process. The model maintains high fidelity across both major mandatory and non-mandatory activities, with average JSD values consistently below 0.06. Notably, our generated patterns and GPS data capture early morning activities (before 6 AM) that are absent in HTS data, addressing known survey limitations such as underreporting and rounding effects, for example, people typically fail to report some emergency travel activities and tend to make the pattern more routine (Bricka et al., 2012). This demonstrates our framework’s ability to balance GPS and survey data characteristics while maintaining demographic representativeness.
6.3 Spatial Validation
We utilized the generated output from the prior pipeline to assign locations in the context of LA County using Algorithm 3, with a spatial division of eight sub-regions (Figure 13(a)). The model was first calibrated using 100,000 agents from the SCAG Activity-Based Model dataset, then scaled to generate mobility patterns for 1 million agents.
To ensure the reliability of our validation results, we implemented a rigorous three-stage calibration process following the methodology established in He et al. (2024). First, we calibrated the MATSim-based transportation network using an iterative proportional adjustment algorithm that minimizes the sum of squared differences between simulated and observed traffic volumes across 100+ Caltrans PeMS detector stations on the top 10 highest-traffic freeways in LA County. The calibration algorithm iteratively adjusts link capacity factors starting from initial values of 0.8, continuing until convergence criteria are met (typically 6 iterations achieving <5% relative difference across time periods). Second, we validated spatial trip distributions against SCAG ABM benchmarks, achieving cosine similarity of 0.997 for origin-destination matrices. Third, we conducted temporal validation using real-world traffic patterns, confirming realistic peak-hour dynamics and congestion onset/dissipation patterns.
@@ 6.3 Spatial Validation
The spatial distribution of activities across sub-regions demonstrates the effectiveness of our location assignment method (Figure 13(a)). The high cosine similarity (0.997) between the origin-destination matrices of our model and SCAG ABM validates the preservation of regional flow patterns.
@@
Traffic simulation on the LA transportation network (Figure 13(b)) reveals strong system-level performance. The hourly VMT and traffic speed patterns closely match the benchmark SCAG ABM results, with MAPEs of 4.97 and 1.16 respectively (Figure 14(b)). These metrics indicate accurate reproduction of aggregate travel patterns.
@@
We conducted detailed analysis on a major segment of Interstate 405, comparing our results with both SCAG ABM predictions and PeMS real-world observations. The analysis reveals that directional variations in congestion patterns are well-captured. For northbound traffic, we observed a MAPE of 5.85 for volume and 4.45 for speed. Similarly, for southbound traffic, we measured a MAPE of 9.32 for volume and 4.36 for speed. The corridor-level results (Figure 15) demonstrate the model’s capability to reproduce fine-grained traffic dynamics, including peak-hour congestion and directional flow patterns.
7 Egypt Case Study
7.1 Regional Adaptation in Data-Scarce Contexts
@@
7.2 Transfer Learning and Model Training
@@
Iteration_1 0.0055 ± 0.0009 0.0052 ± 0.0004 0.0027 ± 0.0005
Iteration_2 0.0038 ± 0.0006 0.0035 ± 0.0003 0.0021 ± 0.0004
Iteration_3 0.0039 ± 0.0004 0.0025 ± 0.0002 0.0021 ± 0.0003
Iteration_4 0.0027 ± 0.0002 0.0034 ± 0.0003 0.0029 ± 0.0004
Iteration_5 0.0043 ± 0.0005 0.0042 ± 0.0004 0.0045 ± 0.0006
Iteration_6 0.0029 ± 0.0003 0.0049 ± 0.0006 0.0046 ± 0.0005
To adapt the base model trained on the NHTS dataset, we employed transfer learning to the Egypt case scenario, leveraging the semi-supervised iterative refinement process 1. The adaptation process retained universal temporal patterns while progressively aligning with the distinct cultural and regional mobility characteristics of Egypt. We then applied the progressive freezing method during the training process, as mentioned in the methodology. Initially, only the MLP and embedding layers were fine-tuned, allowing for rapid adaptation of activity classifications and temporal representations. In subsequent iterations, the transformer layers were gradually unfrozen to refine the model’s understanding of intermediate and complex activity dependencies. Its evolution through the semi-supervised transfer learning using Egypt GPS data is visualized and quantified in Figure 16 and Table 7.
@@
The Egypt TUS data was aggregated to derive target activity duration distributions across the six demographic groups. Adapter training followed the same computational configuration as the case of LA County Case Study
The inference process uses the same setup as the prior case study, generating synthetic activity chains that closely aligned with the target distributions. As shown in Table 8, the JSD values for all demographic groups were low (ranging from 0.0054 for mid-aged males to 0.0302 for mid-aged females), indicating strong alignment between the generated and target distributions. This underscores the framework’s ability to replicate realistic activity patterns even in a data-scarce context.
@@
Young Male (0.0067) 842.2/790.6 95.0/90.7 347.6/308.2 4.2/15.8 15.0/1.4 33.5/38.9 43.0/31.7 32.4/56.2
Young Female (0.0172) 786.1/761.8 76.8/70.6 316.6/313.9 3.1/53.3 14.4/4.3 131.5/148.3 1.4/8.6 36.2/53.3
Mid-Aged Male (0.0054) 684.1/688.3 342.4/348.5 12.7/10.1 8.9/14.4 15.0/15.8 136.7/126.7 4.1/4.3 119.9/112.3
Mid-Aged Female (0.0302) 881.7/832.3 199.1/146.9 63.7/0.0 8.1/24.5 19.7/5.8 146.7/126.7 0.3/4.3 36.2/53.3
Elderly Male (0.0141) 752.7/764.6 80.2/56.2 2.2/2.9 4.6/18.7 17.4/47.5 186.9/197.3 1.3/5.8 135.4/108.0
Elderly Female (0.0198) 848.6/853.9 107.9/79.2 0.0/17.3 3.1/18.7 10.9/4.3 164.3/142.6 0.0/11.5 46.6/106.6
@@
7.3 Observations on Generated Patterns
The generated activity chains demonstrated high consistency with target distributions, particularly for mandatory activities. Mid-aged males exhibited significantly longer work durations compared to mid-aged females, reflecting the disparity in employment rates (65% for males vs. 18% for females, as reported by the UN). Work durations for females in Egypt were notably lower than those in Los Angeles, illustrating the framework’s sensitivity to cultural and socio-economic differences. Religious activities are prominent across all demographic groups, with peaks reflecting the significance of prayer times in Egyptian society. Recreational activities are observed to be less frequent, consistent with the conservative cultural context of Egypt compared to Western regions like Los Angeles.